OpenStudy (anonymous):

How many units apart are the points (8, 11) and (-1, -1)?

OpenStudy (lgbasallote):

use distance formula. do you know what the formula is?

OpenStudy (anonymous):

Isn't it the Pythagorean theorem? I know that, I just don't know how to use it.

OpenStudy (anonymous):

it is pythagoras exactly

OpenStudy (lgbasallote):

pythagorean theorem? sort of \[\sqrt{(x_2 - x_1)^2 + (y_2 - y_1)^2}\] and it's pythagoras?

OpenStudy (anonymous):

from -1 to 8 is 9 units from -1 to 11 is 12 units. so \[d^2=9^2+12^2\] and therefore \[d=\sqrt{9^2+12^2}\]

OpenStudy (anonymous):



OpenStudy (anonymous):

Is it 225?

OpenStudy (anonymous):

Using Pythagoras

OpenStudy (anonymous):

It's 15*